OpenStudy (anonymous):

Please help! I need help with Linear Functions! I will fan and medal!!! f(x) = -3x + 2 Linear or Nonlinear? m = b =

OpenStudy (anonymous):

it is linearly because it looks like \[f(x)=mx+b\]

OpenStudy (anonymous):

so how do I figure out what m is and what b is?

OpenStudy (anonymous):

\[\large f(x) = \color{blue}mx + \color{red}2 \]\[\large f(x) = \color{blue}{-3}x + \color{red}2 \]

OpenStudy (anonymous):

so x = 5?

OpenStudy (anonymous):

no x is a variable

OpenStudy (anonymous):

also, i don't see any five in the question

OpenStudy (anonymous):

lets try again

OpenStudy (anonymous):

I'm sorry and ok

OpenStudy (anonymous):

\[\large f(x) = \color{blue}mx + \color{red}b\]\[\large f(x) = \color{blue}{-3}x + \color{red}2\]

OpenStudy (anonymous):

match up the colors

OpenStudy (anonymous):

\[\huge\color{blue}m=\color{blue}{-3}\]

OpenStudy (anonymous):

Oh I see!

OpenStudy (anonymous):

so if m= -3 then b=2?

OpenStudy (anonymous):

and \[\huge \color{red}b=\color{red}2\] yes

OpenStudy (anonymous):

I get it now! Thank you!

OpenStudy (anonymous):

will you help me with some more?

OpenStudy (anonymous):

f(x) = x2 - 1 Linear or Nonlinear? m = b =

OpenStudy (anonymous):

I know that this one is non linear because it has an exponent

OpenStudy (anonymous):

so how do I know what m or b is?

OpenStudy (anonymous):

the graph is  so how can it be linear ^^

OpenStudy (loser66):

not linear--> no m or b DAT SIT

OpenStudy (anonymous):

linear equation is the one which gives straight line ~ its basic remember

OpenStudy (loser66):